OpenStudy (anonymous):

A right circular cylinder has a base whose area is 154 square feet. If the height of the cylinder is 19 feet, what is its lateral surface area? Use π =

OpenStudy (anonymous):

Solve the following for r:\[154=\pi r^2 \]\[r=\sqrt{\frac{154}{\pi }} \]Lateral area is the circle's circumfrence mulitplied by the cylinder's height.\[\left(\pi 2 \sqrt{\frac{154}{\pi }}\right)19=38 \sqrt{154 \pi }=835.832 \text{ ft}^2 \]